OpenStudy (anonymous):

Help ! Solve: 2(x - 3) + 2(x - 7) = 40 A. x = 12.5 B. x = 15 C. x = 5 D. no solution E. infinitely many solutions

OpenStudy (anonymous):

2x-6+2x-14=40 4x-20=40 x=60/4 x=15

OpenStudy (bahrom7893):

order is right